• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 19:56:03 on February 26, 2009

    Entry number 263559

    keyword=Right Q3

    During the access earlier today we found the Current Lead Monitor wires on the flags on the magnet were reversed. Heidi corrected this. The supply seems to behave correctly, now. We did not ramp it to full field. Also, I inadvertently left the output voltage limit set to 1.20V. So, for now just run it where you need it. We will fix the voltage limit and ramp to 1600A at the next opportunity.